Grade 12Physical Chemistry

Elevation in bp of molar solution of glucose with density 1.2 is

Aniket
9 Years agoGrade 12

1 Answer

varun
9 Years ago
 HELLO ANIKET,
density = 1.2g/ml
(assume we have 1 litre solution)
so molarity = 1M
total mass of solution = 16 X density X 1000ml/L
                                  = 1200gm of solution
now,
1M of glucose => 1X180 = 180gm of glucose
weight of the solvent = 1200-180 = 1020gm 
                                                    = 1.02kg
now,
Molality = 1 mol of solute/1.02 kg of solvent
              =1/1.02
                = 0.98
so,
Tb = m X Kb
     0.98Kb
THANK YOU